### v1

# Asymmetric Bias in Text-to-Image Generation with Adversarial Attacks

Abstract
The widespread use of Text-to-Image (T2I) models in content generation requires careful examination of their safety, including their robustness to adversarial attacks. Despite extensive research into this, the reasons for their effectiveness are underexplored. This paper presents an empirical study on adversarial attacks against T2I models, focusing on analyzing factors associated with attack success rates (ASRs). We introduce a new attack objective - entity swapping using adversarial suffixes and two gradient-based attack algorithms. Human and automatic evaluations reveal the asymmetric nature of ASRs on entity swap: for example, it is easier to replace “human” with “robot” in the prompt “a human dancing in the rain.” with an adversarial suffix, but is significantly harder in reverse. We further propose probing metrics to establish indicative signals from the model’s beliefs to the adversarial ASR. We identify conditions resulting in a 60% success probability for adversarial attacks and others where this likelihood drops below 5%.
1 Introduction
Figure 1: Overview of new attack objective, its asymmetric success rate, and the underlying cause of said asymmetry.
The capabilities of Text-to-Image (T2I) generation models, such as DALL-E 2 (Ramesh et al., 2022), DALL-E 3 (Betker et al., 2023), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022), have improved drastically and reached commercial viability. As with any consumer-facing AI solution, the safety and robustness of these models remain pressing concerns that require scrutiny.
The majority of research related to T2I safety is associated with the generation of Not-Safe-For-Work (NSFW) images with violence or nudity (Qu et al., 2023; Rando et al., 2022; Tsai et al., 2023). To counter this, pre-filters that check for NSFW texts and post-filters that check for NSFW images are used Safety-checker (2022). These filters, however, are not infallible (Rando et al., 2022), and research into bypassing them with attacks, termed ‘jailbreaking’ is advancing (Yang et al., 2023b, a; Noever and Noever, 2021; Fort, 2023; Galindo and Faria, ; Maus et al., 2023; Zhuang et al., 2023). These attacks typically view the shift from normal to NSFW-triggering adversarial prompts as a singular challenge, without sufficiently investigating the reasons behind these attacks’ effectiveness.
On the other hand, few explainability studies have examined Text-to-Image (T2I) models’ capabilities and vulnerabilities, often analyzing model features without adversarial attacks. The findings include the models’ tendency to generate better composition-aware content  (Kong et al., 2023) but without understanding it (West et al., 2023) and the existence of compositional distractors (Hsieh et al., 2023). We identify a specific bias of T2I models linked to the adversarial attack success rates, bridging the gap between the attack and explainability research. We demonstrate the asymmetric bias of T2I models by conducting adversarial attacks in a novel entity swapping scenario, in contrast to the existing setup of removing objects Zhuang et al. (2023) or inducing NSFW content Yang et al. (2023b, a). This setup enables us to investigate the attack success rate in a cyclical setting.
To study the adversarial attack in a cyclic manner, the attack itself needs to be potent with a high attack success rate. We found that directly applying gradient-based attacks like AutoPrompt Shin et al. (2020) or LLM attack Zou et al. (2023) does not guarantee a high attack success rate on T2I models due to being unnecessarily conservative in their approaches to generating adversarial perturbations and the larger vocabulary size of T2I text-encoder Radford et al. (2021) compared to LLMs Touvron et al. (2023); Chiang et al. (2023) (${}50K$ vs. $32K$) which leads to a larger unrestricted search space ($\sim 10^{24}$ vs. $\sim 10^{23}$ for 5 token suffixes).
Our novel setup [^1] Experimental results and code at https://github.com/Patchwork53/AsymmetricAttack and efficient adversarial attack have enabled us to observe an asymmetric attack success rate associated with entity swap. Initially, we hypothesized that long-tail prompts with high perplexity would be more vulnerable to attacks. Surprisingly, we found no strong correlation between the Attack Success Rate (ASR) and the perplexity of the prompt. However, with our proposed measure that evaluates the internal beliefs of CLIP models, we detected indicative signals for ASR, which help identify examples or prompts that are more susceptible to being attacked. Our contributions can be summarized as follows:
  1. 1.
    We introduce a new attack objective: replacing entities of the prompt using an adversarial suffix, allowing us to study whether the attack success rate is associated with the underlying data distribution of prompts. We also propose an automated evaluation with strong human correlation for measuring entity swapping attack success rate (Figure 0(a)).
  2. 2.
    We propose a robust and efficient query-free targeted attack method, utilizing a gradient-based search algorithm to effectively identify adversarial suffixes. This powerful attack method reveals a clear distinction in the attack success rate when swapping two entities in opposite directions, indicating an asymmetry in adversarial attacks (Figure 0(b)).
  3. 3.
    We propose a new measure that is tied to the asymmetric bias of T2I models. This helps us identify vulnerable preconditions, allowing us to estimate the likelihood of a successful attack without actually performing the attack (Figure 0(c)).
2 Related Works
Adversarial Attacks
Adversarial attacks, which perturb inputs to cause models to behave unpredictably, have been a long-studied area in the field of adversarial robustness (Szegedy et al., 2013; Shafahi et al., 2018; Shayegani et al., 2023b). Earlier studies on adversarial attacks focused on neural networks such as convolutional neural networks (Athalye et al., 2018; Hendrycks and Dietterich, 2018), while recent work has shifted towards examining large language models (LLMs) (Shin et al., 2020; Zou et al., 2023; Liu et al., 2023c; Mo et al., 2023; Cao et al., 2023), vision language models (VLMs) (Dong et al., 2023; Khare et al., 2023; Shayegani et al., 2023a), and Text-to-Image (T2I) models detailed below.
Attacks on T2I Models
Zhuang et al. (2023) were among the first to demonstrate that a mere five-character perturbation could significantly alter the images generated. Tsai et al. (2023) and SneakyPrompt (Yang et al., 2023b) each proposed an adversarial attack using genetic algorithms and reinforcement learning algorithms, respectively, to perturb safe prompts into generating NSFW content. VLAttack (Yin et al., 2023), MMA-Diffusion (Yang et al., 2023a), and INSTRUCTTA (Wang et al., 2023) demonstrated that cross-modality attacks can achieve higher success rates than text-only attacks. For defense, Zhang et al. (2023) proposed Adversarial Prompt Tuning (AdvPT) to enhance the adversarial robustness of the image encoder in T2I models.
Vulnerability Analysis
Previous studies (Ilyas et al., 2019; Shafahi et al., 2018; Brown et al., 2017) have explored reasons for model vulnerability to adversarial attacks, especially in image classification. Ilyas et al. (2019) suggested that adversarial examples stem from non-robust features in models’ representations, which are highly predictive yet imperceptible to humans. However, non-robust features for predictive models might not fully explain the effectiveness of adversarial attacks against generative models. Subhash et al. (2023) suggested that adversarial attacks on LLMs may act like optimized embedding vectors, targeting semantic regions that encode undesirable behaviors during the generation process.
Our study, distinct from prior research, focuses on analyzing factors in the model’s beliefs that are associated with attack success rates, introducing a unique entity-swapping attack setup. We developed a discrete token-searching algorithm for targeted attacks and identified asymmetric biases in attack success rates linked to the model’s internal bias. Our experiments emphasize the interplay between prompt distributions, model biases, and attack success rates.
3 Entity Swapping Attack
This section describes the proposed setup of the entity-swapping attack and the corresponding evaluation metric. Designing a new attack scenario may be straightforward, but developing a suitable measure is non-trivial. We also find that, unlike the ease of triggering NSFW image generation, the current gradient-based adversarial suffix search algorithm struggles with targeted entity-swapping attacks. Towards this end, we propose two efficient discrete token search algorithms for the attack, resulting in improved success rates in entity-swapping attacks.
3.1 T2I Model Setup
We study entity-swapping attacks using Stable Diffusion (Rombach et al., 2022), a widely-used T2I model based on a denoising diffusion probabilistic model with a U-Net architecture. It uses cross-attention and CLIP (Radford et al., 2021) for text-image alignment, and a variational autoencoder (Kingma and Welling, 2013) for latent space encoding.The model’s dependence on CLIP text embeddings increases its vulnerability to adversarial attacks. See Appendix C for details.
The CLIP text-encoder transforms prompt tokens $x_{1:n}$ into hidden states $h_{1:n}$ with dimension $D$, which the U-Net component in Stable Diffusion uses to attend via cross-attention. The operation $\mathcal{H}$ represents the combined process of encoding tokens $x_{1:n}$ and reshaping the output hidden states into a one-dimensional vector of shape $n\times D$.
$\mathcal{H}(x_{1:n})=\text{Flatten}(\text{CLIP}(x_{1:n}))$ (1)
We pad initial and target prompts to 77 tokens, the maximum CLIP supports, to handle varying token counts for attacks.
Figure 2: Targeted replacement of entities (blue text) using adversarial suffixes (red highlight). Examples generated using Stable Diffusion v2-1-base. This attack setup allows us to study the correlation between prompt distribution and attack success rate.
3.2 Proposed Attack
To study how the underlying data distribution of prompts relates to the success rate of attacks, we designed a simple yet insightful setup for adversarial attacks: targeted entity swapping. We carefully collected 100 plausible prompts and identified a contrasting pair of entities for each to study the entity-swapping attack within the context of adversarial attacks. Our approach is straightforward: rather than attempting to manipulate T2I to produce violent images or to remove a given object without restrictions, we aim to replace an object in the image with another targeted one. This approach also allows us to explore the feasibility of reverse attacks by inserting adversarial tokens. Illustrations and examples of our attack setup can be found in Figure 2.
The attack, designed to launch from the embedding space, aims to maximize a score function that measures the shift from the input tokens’ embeddings $\mathcal{H}(x^{T}_{1:n})$ towards the target tokens’ embeddings $\mathcal{H}(x^{S}_{1:n})$ using cosine similarity:
$\begin{split}\mathcal{S}(x_{1:n})=w_{t}\times\text{cos}(\mathcal{H}(x^{T}_{1:n}),\mathcal{H}(x_{1:n}))-\\ w_{s}\times\text{cos}(\mathcal{H}(x^{S}_{1:n}),\mathcal{H}(x_{1:n}))\end{split}$ (2)
where $w_{t}$ and $w_{s}$ are weighing scalars and cos denotes the standard cosine similarity metric between two one-dimensional text embeddings.
Optimizing $\mathcal{S}$ is challenging due to the discrete token set and the exponential search space ($k^{|V|}$ for $k$ suffix tokens), making simple greedy search intractable. Current solutions, like HotFlip Ebrahimi et al. (2017) and concurrent work applied to Stable Diffusion Yang et al. (2023a), take gradients with respect to one-hot token vectors and replace tokens for all positions in the suffix simultaneously. The linearized approximation of replacing the $i^{th}$ token, $x_{i}$, is computed by evaluating the following gradients:
$\nabla_{e_{x_{i}}}\mathcal{L}(x_{1:n})\in\mathbb{R}^{|V|},\quad\mathcal{L}(x_{1:n})=-\mathcal{S}(x_{1:n})$ (3)
where ${e_{x_{i}}}$ denotes the one-hot vector representing the current value of the $i^{th}$ token.
3.3 Proposed Optimization Algorithms
We propose two efficient algorithms for finding adversarial suffix tokens against Stable Diffusion, adapting the Greedy Coordinate Gradient algorithm (Zou et al., 2023; Shin et al., 2020) with a focus on maximizing $\mathcal{S}$ for semantic similarity rather than minimizing $\mathcal{L}$ for outlier penalization.
Single Token Perturbation.
This is a straightforward modification of the Greedy Coordinate Gradient algorithm (Zou et al., 2023) with the negative log-likelihood function replaced with our loss defined in Eqn. 3. At each optimization step, our algorithm selects $k$ tokens with the highest negative loss as replacement candidates, $\chi_{i}$, for each adversarial suffix position $i$. It then creates $B$ new prompts by randomly replacing one token from the candidates. Each prompt in $B$ differs from the initial prompt by only one token. The element of B with the highest $\mathcal{S}$ is then assigned to $x_{1:n}$. We repeat this process $T$ times.
Multiple Token Perturbation.
Inspired by the effectiveness of Genetic Algorithms (Sivanandam et al., 2008) seen in untargeted T2I attacks (Zhuang et al., 2023), we hypothesize that replacing multiple tokens simultaneously could enhance convergence speed. Unlike the LLMs targeted by Zou et al. (2023) such as LLaMA (Touvron et al., 2023) and Vicuna (Chiang et al., 2023), CLIP models operate more like bag-of-words, fully capturing syntactical and schematic token relations (Yuksekgonul et al., 2022). Practically, we observe that replacing multiple tokens per optimization step yields a higher success rate within the same computational budget.
In detail, the algorithm selects $k$ tokens and creates $B$ new prompts by randomly replacing multiple token positions. Drawing inspiration from the classic exploration versus exploitation strategy in reinforcement learning (Sutton and Barto, 2018), we initially replace all tokens and then gradually decrease the replacement rate to $25\%$. Figure 2 illustrates some adversarial suffixes generated using this algorithm. Details of both algorithms are provided in the Appendix A.
Token Restrictions
Using the two algorithms for adversarial suffix searching, we observed that 5 adversarial tokens often include words semantically similar to the targets. To ensure imperceptible attacks, we can limit the adversarial suffix to a set of tokens, denoted as $\mathcal{A}$. By setting $V-\mathcal{A}$ gradients to infinity before the Top-k operation, we ensure only $\mathcal{A}$ tokens are chosen. This method allows us to mimic QFAttack Zhuang et al. (2023), modifying the image with arbitrary 5-letter suffixes, as depicted in Figure 3, or to exclude target synonyms in the attack suffix.
Figure 3: Our targeted attack can emulate the restricted token attack (untargeted) from Zhuang et al. (2023) using five ASCII tokens with Stable Diffusion 1.4. The blue text indicates the part we want to remove.
3.4 Proposed Attack Evaluation
To assess the success of a targeted entity swapping attack, we use a classifier to verify if the generated image matches the input or target prompt. Given a tuple (input text, target text, generated image), we define a classifier $\mathcal{C}$ as follows:
[…]
We define class 00 as cases where the adversarial token search algorithm gets stuck at a local maximum, failing to reach the target within the compute budget. For example, when trying to change “A backpack in a forest” to “A cabin in a forest”, we noticed that some of the generated images depicted “People in a forest” or “A cabin and a backpack in a forest” instead. Class $+1$ alone indicates a successful attack, but this three-class framework enables a more comprehensive comparison between human judgments and our proposed classifiers.
Attack Success Rate (ASR).
Stable Diffusion supports the generation of multiple images in parallel. We define an adversarial suffix as being successful if the target text is a suitable caption for the majority of images (generated in parallel). For example, when trying to change “A backpack in a forest” to “A cabin in a forest”, if we generate 5 images with adversarial suffix appended prompt “A backpack in a forest titanic tycoon cottages caleb dojo, we will consider the adversarial suffix successful by majority vote if 3 or more images match the target prompt.
Human Evaluations/Labels
We gather evaluations from three human evaluators for 200 random samples by presenting them a WebUI (Appendix E) with the generated image and two checkboxes for input text and target text. They are instructed to select texts that match the image and can select one, both, or neither, i.e. into three classes as established in Eqn. 4.
[…]
Choice of the Classifier.
We generate multiple attack suffixes and images for each input-target pair to empirically determine attack success rates. Due to the large volume of images, we employ human evaluators for a subset and VLM-based classifiers for the full set evaluation. Specifically, we test InstructBLIP Liu et al. (2023a), LLaVA-1.5 Liu et al. (2023b), and CLIP Radford et al. (2021), and compare their performance with human labels.
Model # Classes Accuracy F1
InstructBLIP 3 0.79 0.75
LLaVA-1.5 3 0.76 0.74
CLIP 3 0.62 0.55
CLIP-336 3 0.60 0.55
InstructBLIP 2 0.86 0.84
LLaVA-1.5 2 0.83 0.81
CLIP 2 0.70 0.69
CLIP-336 2 0.68 0.67
[…]
Figure 4: Comparison of pair-wise attack success rate on 50 pairs using Multiple Token Perturbation Algorithm.
For auto-regressive VLMs like InstructBLIP and LLaVA-1.5, we use the prompt ‘Does the image match the caption [PROMPT]? Yes or No?’, replacing [PROMPT] with input or target text for image evaluation. For CLIP models, an image is classed as $+1$ if its target text similarity is above $1-\gamma$ and its input text similarity is below $\gamma$, and $-1$ for the reverse case. All other cases are classified as 00.
Table 1 indicates that InstructBLIP outperforms other models such asLLaVA-1.5, CLIP-ViT-L14, and CLIP-ViT-L14-336. Given that CLIP models produce continuous values, we present their performance using the optimal threshold $\gamma$ ($\gamma_{CLIP}=0.0034$ and $\gamma_{CLIP-336}=0.0341$)that maximizes the F1 score against human labels. Consequently, we proceed with using InstructBLIP as the evaluation classifier in subsequent sections.
4 Experiments and Results
This section presents the experimental details and results of adversarial attacks for entity swapping setups, involving the insertion of adversarial suffixes.
4.1 Experimental Setups
We run all experiments, including the attacks, image generation, and VLM evaluation, on a single NVIDIA 3060 12GB and 16GB RAM. We use Stable Diffusion v2-1-base for all experiments unless stated otherwise.
We evaluate 100 input-target pairs to compare the effectiveness of Single and Multiple Token Perturbation. We run each algorithm 10 times per pair with $T=100$ steps with $k=5$ and $B=512$, which yields 10 adversarial attacks per pair. For each attack, we generate 5 images. The two algorithms are evaluated against each other on $100\times 10\times 5=5000$ generated images. We set $w_{t}=w_{s}=1$ in Eqn. 2 for the experiments.
4.2 Overall Attack Results
Using the same hyperparameters and compute budget, our Multiple Token Perturbation algorithm outperforms the Single Token Perturbation used in LLM attacks, with ASR $26.4\%$ and $24.4\%$ respectively over 100 input-target pairs on 5000 generated images. Zou et al. (2023) showed that Single Token Perturbation was an effective adversarial suffix-finding strategy for LLMs. However, the CLIP text is relatively lightweight compared to LLMs and behaves more like a bag-of-words model Yuksekgonul et al. (2022). We find that updating multiple tokens at each time step leads to faster convergence, likely because CLIP demonstrates a reduced emphasis on the semantic and syntactical relationships between tokens, in comparison to LLMs. Our findings corroborate the effectiveness of the Genetic Algorithm in Zhuang et al. (2023), which resembles multiple token perturbations but operates in an untargeted setting without a gradient-based algorithm. However, it’s important to note that Single Token Perturbation could potentially achieve the same ASR as Multiple Token Perturbation if not limited by compute resources. We employ Multiple Token Perturbation for all subsequent experiments analyzing asymmetric attacks.
4.3 Forward and Backward Attack Results
One of our key findings is the strong asymmetry of adversarial attack success rate, as illustrated in Figure 4. For instance, attacks from ‘A swan swimming in a lake.’ to ‘A horse swimming in a lake.’ failed in all ten attempts, whereas the reverse direction achieved an ASR of 0.9. Such asymmetric ASRs are common; however, in other cases, the forward and backward ASRs aren’t inversely proportional. For example, both directions between ‘A man reading a book in a library.’ and ‘A woman reading a book in a library.’ have moderate ASRs of 0.7 and 0.5, respectively, while pairs like (‘A dragon and a treasure chest.’, ‘A knight and a treasure chest.’) fail in both directions. Inspired by these asymmetric observations, we further conducted experiments to analyze the relationship between prompt distribution and attack success rate.
5 Asymmetric ASR Analysis
This section discusses our experiments to analyze the asymmetric ASR observed in Section 4.3. We aim to investigate the model’s internal beliefs that may lead to these distinct attack success rate (ASR) differences from opposite directions. We propose three potential factors for this asymmetry: the difficulty of generating the target text (BSR, Eqn. 5), the ‘naturalness’ of the target text relative to the input text ($\Delta_{1}$, Eqn. 6), and the difference in distance from the target text to the baseline compared to that from the input text ($\Delta_{2}$, Eqn. 7).
5.1 Probe Metrics
We initially speculated that the ASR might be related to the difficulty of generating the target prompt, leading us to evaluate the Base Success Rate (BSR) of target generation.
BSR=Successful GenerationsGeneration $\text{BSR}=\frac{\text{Successful Generations}}{\text{Generation Attempts}}$ (5)
BSR assesses T2I model’s likelihood of producing a target from a PAD token, for example, answering what would Stable Diffusion generate if prompted with ”A _____ swimming in a lake”. Figure 5 shows that T2I models have inherent biases towards certain objects, possibly due to biases in its training data. T2I models may struggle to generate images absent from their training data, as evidenced by our experiments where even simple scenes like “A dragon guarding a treasure” are inconsistently produced. Therefore, if T2I models struggle with the target alone, adversarial attacks aimed at generating them are likely to be even more challenging.
Figure 5: Base success rate: Images generated by Stable Diffusion (SD) 2.1 when prompted with the PAD token (”!” for SD 2.1).
(a) ASR vs. Perplexity Difference ($\Delta_{1}$ in Eqn. 6)
(b) ASR vs. BSR (of target text)
(c) ASR vs. Baseline Distance Difference ($\Delta_{2}$ in Eqn. 7)
(d) ASR for Negative and Positive $\Delta_{2}$
Figure 6: Correlation of ASR on $\Delta_{1}$, $\Delta_{2}$ and BSR. We find that the Perplexity Difference $\Delta_{1}$ does not correlate with ASR. BSR shows a weak positive correlation and Baseline Distance Difference $\Delta_{2}$ shows a moderate negative correlation with ASR.
We also speculated that the difference in Perplexity $\Delta_{1}$, measuring how natural or plausible a prompt is, might be associated with asymmetric ASR. For instance, “A swan swimming in a lake” is a more natural scene than “A horse swimming in a lake”. We selected the Perplexity (PPL) metric from text-davinci-003 by OpenAI Brown et al. (2020), as it most closely matches human judgment compared to T5 Raffel et al. (2020) and LLaMA 2 Touvron et al. (2023) in our experiments. Perplexity for a sequence $x_{1:n}$ is defined as $\text{PPL}(x_{1:n})=e^{-\frac{1}{n}\sum_{i=1}^{n}\log P(x_{i}|x_{1:i-1})}$, and we define the difference, noted as $\Delta_{1}$:
$\mathbf{\Delta_{1}}(x^{T}_{1:n},x^{S}_{1:n})=\text{PPL}(x^{T}_{1:n})-\text{PPL}(x^{S}_{1:n}).$ (6)
Additionally, we introduce a new metric termed Baseline Distance Difference, denoted as $\Delta_{2}$. This combines insights from the probe metric BSR and $\Delta_{1}$, but measures distances using cosine similarity of CLIP embeddings instead of LLM’s perplexity.
$\begin{split}\mathbf{\Delta_{2}}(x^{T}_{1:n},x^{S}_{1:n})=\text{cos}(\mathcal{H}(x^{T}_{1:n}),\mathcal{H}(x^{B}_{1:n}))\\ -\text{cos}(\mathcal{H}(x^{S}_{1:n}),\mathcal{H}(x^{B}_{1:n})).\end{split}$ (7)
5.2 Results
We evaluate 50 random prompt pairs in both directions, resulting in $50\times 2\times 10\times 5=5000$ generated images. To find the base success rate defined in Eqn. 5, we generate 64 images for each of the 50 random prompts and count the number of generations (out of 64) that were successful.
We find that $\Delta_{1}$ has a negligible correlation with ASR (Pearson $r=0.05$ and Spearman $\rho=-0.06$, see in Figure 5(a)). This is counter-intuitive because one might expect a target with lower perplexity compared to the input text should be easier to generate through an adversarial attack.
We also observed that ASR has a weak positive correlation with BSR (see Figure 5(b)) and a moderate one with $\Delta_{2}$ (see Figure 5(c)). In particular, Figure 5(d) shows the mean ASR when $\Delta_{2}<0$ is 0.40, whereas it drops to just 0.12 when $\Delta_{2}\geq 0$. Thus, $\Delta_{2}$ enables us to estimate, to some extent, the likelihood of a successful adversarial attack.
5.3 Predictor for Successful Attack
Considering the observed reasonable correlations of BSR (of the target text) and $\Delta_{2}$ with attack success rates, this section explores whether combining these two indicators can predict the likelihood of a successful entity-swapping attack.
BSR $\boldsymbol{\Delta_{2}}$ Samples Avg. ASR
$<0.9$ $<0$ 23 0.174
$<0.9$ $\geq 0$ 19 0.047
$\boldsymbol{\geq 0.9}$ $\boldsymbol{<0}$ 27 0.6
$\geq 0.9$ $\geq 0$ 31 0.172
Table 2: Average ASR for different combinations of BSR and $\Delta_{2}$. From the holdout set performance, we set $0.9$ as the threshold to classify BSR as high.
Table 2 demonstrates that our probe metric provides a reliable success predictor for attacks: when BSR (of the target text) is high and $\Delta_{2}$ is negative for a given input-target text pair, adversarial attacks have a 60% chance of success, dropping to about 5% when BSR is low and $\Delta_{2}$ is positive. Considering both BSR and $\Delta_{2}$ together improves the prediction of an attack’s success likelihood. However, the need for BSR precomputation is a limitation, and finding ways to approximate BSR without precomputation is an area for future research.
We also identified factors akin to existing research on general elements associated with attack success rates, like the length of the adversarial suffix. These factors, together with our experimental results, are detailed in Appendix D.
6 Conclusion
This paper presents an empirical study on adversarial attacks targeting text-to-image (T2I) generation models, with a specific focus on Stable Diffusion. We define a new attack objective—entity swapping—and introduce two gradient-based algorithms to implement the attack. Our research has identified key factors for successful attacks, revealing the asymmetric nature of attack success rates for forward and backward attacks in entity swapping. Furthermore, we propose probing metrics to associate the asymmetric attack success rate with the asymmetric bias within the T2I model’s internal beliefs, thus establishing a link between a model’s bias and its robustness against adversarial attacks.
Appendix A Algorithms
Algorithm 1 Single Token Perturbation
0:  Initial prompt $x_{1:n}$, modifiable subset $I$, iterations $T$, loss $\mathcal{L}$, score $\mathcal{S}$, batch size $B$, $k$
1:  for $t\in T$ do
2:     for $i\in I$ do
3:        $\chi_{i}\leftarrow\text{Top-}k(-\nabla_{x_{i}}\mathcal{L}(x_{1:n}))$Compute top-$k$ promising token substitutions}
4:     end for
5:     for $b=1,\ldots,B$ do
6:        $x_{1:n}^{(b)}\leftarrow x_{1:n}$Initialize element of batch}
7:        $x_{i}^{(b)}\leftarrow\text{Uniform}(\chi_{i})$, where $i\leftarrow\text{Uniform}(I)$Select random replacement token}
8:     end for
9:     $x_{1:n}\leftarrow x_{1:n}^{(b^{*})}$, where $b^{*}=\arg\max_{b}\mathcal{S}(x_{1:n}^{(b)})$Compute best replacement}
10:  end for
10:  Optimized prompt $x_{1:n}$
Algorithm 2 Multiple Token Perturbation
0:  Input: Initial prompt $x_{1:n}$, modifiable subset $I$, iterations $T$, loss $\mathcal{L}$, score $\mathcal{S}$, batch size $B$, $k$, $\epsilon_{f}$, $\epsilon_{s}$
1:  $\epsilon\leftarrow\epsilon_{s}$
2:  for $t\in T$ do
3:     for $i\in I$ do
4:        $\chi_{i}\leftarrow\text{Top-}k(-\nabla_{x_{i}}\mathcal{L}(x_{1:n}))$Compute top-$k$ promising token substitutions}
5:     end for
6:     for $b=1,\ldots,B$ do
7:        $x_{1:n}^{(b)}\leftarrow x_{1:n}$Initialize element of batch}
8:        for $i\in I$ do
9:           if $\mathcal{P}(\epsilon)$ then
10:              $x_{i}^{(b)}\leftarrow\text{Uniform}(\chi_{i})$Select random replacement token}
11:           end if
12:        end for
13:     end for
14:     $x_{1:n}\leftarrow x_{1:n}^{(b^{*})}$, where $b^{*}=\arg\max_{b}\mathcal{S}(x_{1:n}^{(b)})$Compute best replacement}
15:     $\epsilon\leftarrow\max(\epsilon_{f},\epsilon_{s}-\frac{t}{T})$Reduce the replacement probability}
16:  end for
16:  Output: Optimized prompt $x_{1:n}$
Appendix B Changing the Number of Adversarial Tokens k
(a) Reducing the number of attack tokens ”a red panda/car in a forest.”. Only displaying the adversarial attack suffixes. 2 tokens are sufficient. ”a red panda in a forest. chained sportscar porsche generates ”a car in a forest”.
(b) Reducing the number of attack tokens ”a guitar/piano in a music store.”. Only displaying the adversarial attack suffixes. All 5 tokens are necessary. ”a guitar in a music store. serendipity upright three bank piano generates ”a piano in a music store.”
Figure 7: Reducing the number of tokens in adversarial prompts. Highly dependent on the starting-target text pair.
Appendix C T2I Model Basics
Stable Diffusion Rombach et al. (2022) is built on a denoising diffusion probabilistic model (DDPM) Ho et al. (2020) framework, utilizing a U-Net architecture for its core operations. Key to its text-to-image capabilities is the cross-attention mechanism, which aligns textual inputs with relevant visual features. Specifically, the U-Net attends to image-aligned text embeddings produced by a CLIP Radford et al. (2021) model. Stable Diffusion also incorporates a Variational Autoencoder Kingma and Welling (2013) to efficiently encode images into a latent space, significantly reducing computational requirements while maintaining image quality. Since text embedding generation using a CLIP model is the first stage in the Stable Diffusion pipeline, it is particularly susceptible to adversarial attacks (Galindo and Faria, ; Zhuang et al., 2023). If an adversary can perturb the text embeddings, later stages in the Stable Diffusion pipeline will reflect the perturbed embeddings.
C.1 Exploiting CLIP’s Embedding Space
The CLIP text-encoder maps the textual prompt tokens $x_{1:n}$, with $x_{i}\epsilon\{1,...,V\}$ where V denotes the vocabulary size, namely, the number of tokens to $x_{1:n}$, where $h_{i}$ is the hidden state corresponding to the token $x_{i}$. The U-Net component in Stable Diffusion attends to all $h_{1:n}$ embeddings using cross-attention. $x_{1:n}$ can be flattened into $\mathbb{\Phi}$, a one-dimensional vector of shape $n\times D$, where D is the embedding dimension (typically 768 for CLIP and its variants). For simplicity, we refer to $\mathbb{\Phi}$ as the text embedding of $x_{1:n}$ from here on. Let $\mathcal{H}$ represent the combined operation for encoding tokens $x_{1:n}$ and reshaping the output hidden states.
$\mathbb{\Phi}=\mathcal{H}(x_{1:n})=\text{Flatten}(\text{CLIP}(x_{1:n}))$ (8)
Since input text and target text can vary in the number of tokens and to allow for an arbitrary number of adversarial tokens, we pad all input and targets to 77 tokens each, the maximum number of tokens supported by CLIP.
C.2 Score Function
The cosine similarity metric approximates the effectiveness of appending adversarial tokens at some intermediate optimization step $t$. Moving away from the input tokens’ embedding and gradually towards the target tokens’ embeddings through finding better adversarial tokens can be thought of as maximizing the following score function, similar to the metric in Zhuang et al. (2023).
$\begin{split}\mathcal{S}(x_{1:n})=w_{t}\times\text{cos}(\mathcal{H}(x^{T}_{1:n}),\mathcal{H}(x_{1:n}))-\\ w_{s}\times\text{cos}(\mathcal{H}(x^{S}_{1:n}),\mathcal{H}(x_{1:n}))\end{split}$ (9)
Here $w_{t}$ and $w_{s}$ are weighing scalars and cos denotes the standard cosine similarity metric between two one-dimensional text embeddings. For simplicity, we set $w_{t}=w_{s}=1$ for all experiments.
C.3 Optimization over Discrete Tokens
The main challenge in optimizing $\mathcal{S}$ is that we have to optimize over a discrete set of tokens. Moreover, since the search space is exponential ($k^{|V|}$ for k suffix tokens), a simple greedy search is intractable. However, we can leverage gradients with respect to the one-hot tokens to find a set of promising candidates for replacement at each token position. We use the negated Score Function as the loss function $\mathcal{L}(x_{1:n})=-\mathcal{S}(x_{1:n})$. Maximizing the score is equivalent to minimizing the loss. Since losses are used for top K token selection, the absolute value of the loss doesn’t matter. We can compute the linearized approximation of replacing the $i^{t}h$ token i, $x_{i}$ by evaluating the gradient
$\nabla_{e_{x_{i}}}\mathcal{L}(x_{1:n})\in\mathbb{R}^{|V|}$ (10)
Where ${e_{x_{i}}}$ denotes the one-hot vector representing the current value of the $i^{t}h$ token. Taking gradient with respect to one-hot vectors was pioneered by HotFlip Ebrahimi et al. (2017) and applied on Stable Diffusion by a concurrent work Yang et al. (2023a). Based on this heuristic, we presented two algorithms for finding adversarial suffix tokens against Stable Diffusion.
Appendix D Additional Determinants of Attack Success
These sections discuss the factors beyond the asymmetric properties that are tied to the attack success rate. We have found factors like whether target token synonyms are allowed, attack suffix length and attack POS types are factors indicating the attack’s success. We also found unlike LLM attacks, adversarial suffixies do not transfer across T2I indicating these models might be harder to attack than single modality models.
D.1 Restricted Token Selection
Emulating QFAttack.
We can restrict certain tokens to emulate QFAttack Zhuang et al. (2023) or prevent the exact target word from being selected. We find that QFAttack can be consistently emulated by restricting token selection to tokens corresponding to ASCII characters. We find that such adversarial suffixes can remove concepts (e.g. “a young man” from “a snake and a young man.” or “on a flower” from “a bee sitting on a flower.”) but fail to perform targeted attacks (e.g. changing “a bee sitting on a flower.” to “a bee sitting on a leaf.”). We suspect that this is mainly because ASCII tokens can perturb CLIP’s embedding but are unable to add additional information to it.
Blocking Selection of Target Tokens.
Another potential use case is preventing the selection of the exact target word. However, we find that the algorithm simply finds a synonym or sub-word tokenization for the target word when the exact target word (token) is restricted. For example, when attempting to attack the input text “a backpack on a mountain.” to “a castle on a mountain.”, restricting the token corresponding to “castle” leads to the algorithm including synonyms like “palace”, “chateau”, “fort” or sub-word tokenization like “cast le” or “ca st le” in the adversarial suffix. We find that the effectiveness of the algorithm isn’t affected when the exact target token is restricted and it still finds successful adversarial suffixes using synonyms (when preconditions are met).
Changing the Number of Adversarial Tokens k
We set the number of adversarial tokens to $k=5$ for all experiments. However, we observe that not all input text-target text pairs require $k=5$. “a red panda/car in a forest.” can be attacked with a few as $k=2$, i.e. “a red panda/car in a forest.” while “a guitar/piano in a music store.” required all $k=5$ (see Appendix B). We leave the comprehensive study on the effect of changing the number of tokens as future work.
D.2 Certain Adjectives Resist Adversarial Attacks
We observed that adversarial attacks targeting certain adjectives like color had a very low ASR. For example, swapping out “red” with “blue” in the prompt “a red car in a city road.” failed in all instances. Further challenging examples include “a red/purple backpack on a mountain.” and “a white/black swan on a lake.”. However, other adjectives like “a sapling/towering tree in a forest” or “a roaring/sleeping lion in the Savannah.” had high ASR in at least one direction. We leave further analysis of this phenomenon as future work.
D.3 Adversarial Suffixes Do Not Transfer across T2I Models
[…]
Appendix E Human Evaluation WebUI
Figure 8: UI presented to human evaluators.